if the sp of 10 pens is equal to the cp of 12 pens, find the profit percentage. 

Grade:7

1 Answers

RAJ KUMAR
6 years ago
10×S.P=12×C.P
S.P/C.P=12/10=6/5
this the ratio of S.P and C.P
if S.P=6x C.P=5x
Profit=6x-5x=x
profit%=(x/5x)×100=100/5=20%